### v3

@@ Abstract
While multi-task learning (MTL) has gained significant attention in recent years, its underlying mechanisms remain poorly understood. Recent methods did not yield consistent performance improvements over single task learning (STL) baselines, underscoring the importance of gaining more profound insights about challenges specific to MTL. In our study, we challenge paradigms in MTL in the context of STL: First, the impact of the choice of optimizer has only been mildly investigated in MTL. We show the pivotal role of common STL tools such as the Adam optimizer in MTL empirically in various experiments. To further investigate Adam’s effectiveness, we theoretical derive a partial loss-scale invariance under mild assumptions. Second, the notion of gradient conflicts has often been phrased as a specific problem in MTL. We delve into the role of gradient conflicts in MTL and compare it to STL. For angular gradient alignment we find no evidence that this is a unique problem in MTL. We emphasize differences in gradient magnitude as the main distinguishing factor. Lastly, we compare the transferability of features learned through MTL and STL on common image corruptions, and find light evidence that MTL can lead to superior transferability. Overall, we find surprising similarities between STL and MTL suggesting to consider methods from both fields in a broader context.
@@ 1 Introduction
In a seminal study, Caruana [4] highlights both advantages and challenges in MTL. On the one hand, certain tasks can exhibit a symbiotic relationship, resulting in a mutual performance enhancement when trained together. It was further suggested that features learned in a MTL scenario rely less on incidental correlations and demonstrate improved transferability. On the other hand, conflicts between tasks can arise and decrease the performance when trained jointly, also known as negative transfer.
@@
The current understanding of MTL is still limited and lacks a deeper comprehension of its underlying mechanisms. To address this gap, our study aims to challenge commonly held paradigms, such as the choice of optimizer, the notion of gradient alignment, gradient magnitudes, and transferability of features.
Our contributions are:
@@
  • We examine the presumption of increased robustness on corrupted data as a result of MTL [24, 37]. We find light evidence that an increased number of tasks can lead to an improved transferability.
Overall, we provide a vast set of experiments and theoretical insights which challenge common paradigms and contribute to a more comprehensive understanding of MTL in computer vision to guide future research.
@@ 2 Related Work
Network architectures focus on the question of how features should be shared across tasks [38, 34, 50, 36]. Multi-task optimization (MTO) aims to resolve imbalances and conflicts of tasks during MTL. Task affinities examine a grouping of tasks that should be learned together to benefit from the joint training [47, 12]. A general overview of recent works in MTL can be found in [48, 43]. Our work focuses on MTO, which we review more thoroughly in the following.
Gradient magnitude methods prevent the dominance of individual tasks by balancing them with task-specific weights. One line of works are loss-weighting methods. Here, weights are determined before any (task-wise) gradient computation and are used for a weighted aggregation of the tasks’ losses. These methods consider either the task uncertainty (UW) [21], rate of change of the losses (DWA, FAMO) [34, 30], the tasks’ difficulty (DTP) [14], validation performance by applying meta-learning (MOML, Auto-$\lambda$) [52, 33], or randomly chosen task weights (RLW) [28]. In line with these, the geometric mean of task losses has been used to handle the different convergence rates of the tasks [8]. An advantage of theses methods is their computational efficiency as the gradient needs to be computed only once for the aggregated loss. Alternatively, other methods consider the task-specific gradients directly, e.g., by normalizing them (GradNorm) [6] or propose a hybrid balancing between task-wise loss and gradient scaling (IMTL, DB-MTL) [32, 27]. Furthermore, there are several adaptions for the multiple-gradient descent algorithm (MGDA) [10], e.g. for applying it efficiently in deep learning setups [44] or by introducing a stochastic gradient correction [11]. Recently, task-wise gradient weights have been estimated by treating MTL as a bargaining problem (Nash-MTL) [40], or considering a stability criterion (Aligned-MTL) [45]. Crucially, in the context of this study, all gradient magnitude methods consider scalar weightings of task-wise gradients within the backbone and/or heads. They do not modify the alignment of task-specific gradient vectors.
@@ 3 Problem Statement
Multi-task learning addresses the problem of learning a set of $T$ tasks simultaneously. We consider a supervised learning setup, use a shared backbone architecture, and learn all tasks together. Formally, given input data $\mathcal{X}$, the goal is to learn a function $f_{\bm{\theta}}(x)$ which maps a point $\mathbf{x}\in\mathcal{X}$ to each task label $y_{t}$ with $t=1,..,T$. The trainable parameters $\bm{\theta}=\{\phi,\psi_{1:T}\}$ consist of shared parameters $\phi$ and task-specific parameters $\psi_{t}$. Training a task $t$ is associated with the loss $\mathcal{L}_{t}(f_{\bm{\theta}}(x);\bm{\theta})$, e.g., a regression or classification loss. We denote respective gradients on the shared and task-specific parameters with $\mathbf{g}^{\phi}_{t}=\nabla_{\phi}\mathcal{L}_{t}$, and $\mathbf{g}^{\psi}_{t}=\nabla_{\psi}\mathcal{L}_{t}$. When training on multiple tasks, the shared parameters $\phi$ needs to be updated w.r.t. all task-wise gradients $\mathbf{g}^{\phi}_{t}$ which requires an appropriate aggregation. A simple solution is to uniformly sum up the task losses $\mathcal{L}=\sum_{t}\mathcal{L}_{t}$ which is referred to as Equal Weighting (EW). However, as tasks might be competing against each other, this can result in negative transfer and thus sub-optimal solutions. One way to deal with this difficulty is to adapt the magnitude of task-specific gradients. This can be achieved by weighting tasks during training, e.g., by scaling different losses $\mathcal{L}=\sum_{t}\alpha_{t}\mathcal{L}_{t}$, where $\alpha_{t}\geq 1$. Note that the $\alpha_{t}$ can change during training. Furthermore, the weighing can also be performed on gradient level to distinguish between shared and task-specific gradients. We refer to those approaches as gradient magnitude methods. Interestingly, the relationship between loss weights, network-updates and learning rate also depends on the optimizer. We show a derivation for SGD and Adam in Section A1.2. Additionally to adapting the gradient magnitude, one can directly adapt the alignment of task-wise gradient vectors within the shared backbone $\tilde{\mathbf{g}}^{\phi}=\mathbf{h}(\mathbf{g_{1}}^{\phi},...,\mathbf{g_{T}}^{\phi})$.
@@ 4 Experiments and results
In this section, we perform several experiments to gain a more profound understanding of multi-task learning (MTL) in computer vision by questioning common paradigms. We compare the impact of Adam and SGD in MTL in Sec. 4.1, examine the process of gradient similarity in different settings in Sec. 4.2, and evaluate the generalization performance on corrupted data in Sec. 4.3. Throughout this evaluation, we repeatedly make use of common setups, which we will specify as follows and in more detail in Appendix A3.
@@ 4.1 Effectiveness of Adam in multi-task learning
Challenged paradigm: The impact of the choice of standard optimizer is often disregarded and varies across studies (see Tab. A1 for an overview) when comparing MTL methods. For instance, Adam [22] was successfully used to show that random/constant weighting of tasks’ losses performs competitive compared to MTO methods [28, 25, 49]. In contrast, many methods proposing adaptive, task-specific weighing methods [32, 21] use stochastic gradient descent with momentum (SGD+mom). In recent works, the optimizer choice converged to Adam and a fixed learning rate schedule [34, 55, 31, 45] without a comparison to SGD+mom.
In this part of our study, we investigate the impact of Adam and SGD+mom in conjunction with common MTO methods. We identify the choice of optimizer as a crucial confounder in the experimental setup. Compared to SGD+mom, we find that the Adam optimizer itself is a quite effective baseline in MTL and to our surprise can be regarded as a loss weighting method from a theoretical viewpoint.
@@ 4.1.1 Toy Task Experiment
Approach: We repeat the experiment of Liu et al. [31] using their original implementation but further test different learning rates and optimizers. They motivate their gradient alignment method CAGrad with a simple toy optimization problem in which their method reliably converges to the minimum of the average loss, while other MTO approaches would either get stuck (e.g., EW) or only convert to any point on the Pareto front (e.g., PCGrad [55], MGDA [44]).
Table 1: Maximum number of iterations after which all seeds in the toy task experiment from CAGrad [31] reach the global minimum for varying MTO method, learning rate, and optimizer combination. In several setups, EW+Adam shows the fastest convergence to the global minimum. If not all seeds converged to the global minimum within 100k iteration steps, we denote it as ’-’. As reported in previous work, we found that PCGrad converges only to some point on the Pareto Front. The best run for each learning rate over all MTO methods is indicated via font type.
@@
Result: For higher learning rates with Adam optimizer, even the simple equal weighting (EW) method reaches the global optimum (cf. Fig. 1, e.g., EW+Adam, lr=0.05) and often converges even faster than dedicated MTO methods such as CAGrad (Tab. 1). Note, original results were shown for learning rate 0.001 using Adam (violett outline) and were, therefore, in favor of CAGrad. Results for additional learning rates are reported in LABEL:tab:cagrad_toytask_niter_full and LABEL:fig:cagrad_toytask_extended.
Conclusion: In summary, the choice of optimizer appears to be more important on the success of the outcome of the experiments than the choice of MTO method, as Adam converges considerably faster and more reliably than GD.
@@ 4.1.2 Experiments on CityScapes and NYUv2
We further extensively test the effectiveness of Adam and its role as a confounder in common MTL datasets for various MTO methods.
Approach: We compare Adam and SGD+mom in combination with any MTO method from equal weighting (EW), uncertainty weighting (UW) [21], random loss weighting (RLW) [28], PCGrad [55], CAGrad [31], IMTL [32] and Aligned-MTL [45], for which we used the implementation from  [29], as well as MTL-IO [41] and AdaTask [51]. We distinguish between any combination of dataset {CityScapes [9], NYUv2 [39]} and network architecture {SegNet [1], DeepLabv3 [5]}. We run experiments for ten different initial learning rates from $[0.5,0.1,0.05,...,0.00001]$ and select the best one w.r.t. to the validation performance. More details are described in Section A3.2. As different models and parameter setups can show preference towards different tasks and metrics, we are interested in those models which are Pareto optimal (PO).
Table 2: Number of Pareto optimal (PO) experiments using either Adam or SGD+momentum as optimizer. Models trained with Adam are consistently more often on the Pareto front compared to those trained with SGD+mom. The number of Adam-based runs that are not dominated by any SGD-based run (PO w.r.t. SGD) is even higher, which does not hold the other way around.
@@
Results: We observe over all experimental setups that Adam performs favorably over SGD+mom (Tab. 2). This especially holds true for experiments on CityScapes where the Pareto front for both network architectures only consists of Adam-based models. Moreover, an even larger number of Adam-based models is not dominated by any model trained with SGD+mom (PO w.r.t. SGD). For NYUv2, Adam still performs stronger but SGD+mom. also occasionally delivers a PO result. For the individual metrics, the predominance of Adam is further visualized in a parallel coordinate plot in Figs. 2 and 4.3. Bold lines indicate the overall Pareto optimal experiments (PO full).
In Appendix A4, we further report best $\Delta_{m}$ results for common MTO methods in combination with Adam or SGD+mom (LABEL:tab:res_cs_segnet, LABEL:tab:res_cs_deeplab, LABEL:tab:res_nyu_segnet and LABEL:tab:res_nyu_deeplab). Again, Adam boosts the overall performance across methods. Furthermore, when comparing the ranking of MTO methods w.r.t. the $\Delta_{m}$ metric, we see that the order can change based on the choice of optimizer, e.g., for Cityscapes with SegNet the best method with Adam is UW but with SGD+mom it is CAGrad. This underlines the importance of the choice of optimizer as a confounder in the experimental setup. Noteworthy, EW with Adam yields Pareto optimal results in three of the four setups (cf. Tab. A2) and is not dominated by any specialized MTO method trained in combination with SGD+mom for all dataset and network combinations. This supports claims questioning the effectiveness of specific MTO methods [25, 49]. Nonetheless, looking at the $\Delta_{m}$-metric and individual metric, we see that sometimes with a small relative performance drop on one metric, significant gains on another metric can be achieved (e.g., Cityscapes+sem.seg. and depth for UW vs EW).
@@ 4.1.3 The reasonable effectiveness of Adam in the context of uncertainty weighting
We show that Adam’s mechanism to estimate a parameter-specific learning rate is partially loss-scale invariant which could contribute to Adams effectiveness in MTL. We demonstrate this partial invariance theoretically and empirically. Furthermore, a full loss-scale invariance can also be shown under mild assumptions for UW [21], which is among the most prevalent loss weighting method in the literature, and related similar variant [27].
The loss-scale invariance of UW can be shown by assuming an optimal solution for the $\sigma$ values similar to [23]. This assumption is mild as this is a 1-dimensional convex optimization problem for each $\sigma$. The invariance can be demonstrated by inserting the analytical solution starting from UW. For example, assuming a Laplacian distribution for simplicity (this can be shown for other distributions as well), we have
@@
where $sg$ is the stop-gradient operator and $c$ is a constant that can be omitted during optimization. Given this equation, we directly see the invariance w.r.t. loss-scalings. For instance, with $\mathcal{L}_{1}\rightarrow\alpha_{1}\mathcal{L}_{1}$ and $\mathcal{L}_{2}\rightarrow\alpha_{2}\mathcal{L}_{2}$, the derivative of the total loss $\mathcal{L}$ remains unchanged. As this invariance is shown on the loss-level, it holds for all gradient updates w.r.t. the head and backbone. Intuitively, this can explain why UW performs strongly in the context of various loss scalings such as measuring depth in centimeters, meters or inches. For further details, we refer to Appendix A1.
@@
Thus, for the network heads, we see a similar effect as for optimal UW that different scalings do not impact the network update. However, this does not hold for the backbone. The full derivation is shown in Appendix A1. We confirm empirically in a handcrafted loss-scaling experiment in Appendix A2 and Figs. A1 and A2 that SGD does not offer any scaling invariance, whereas Adam involves the invariance property for the heads. The optimal UW demonstrates a scaling invariance for the heads and the backbone.
We would like to note that our derivation for Adam is only valid for constant $\alpha_{t}$, e.g., measuring depth in different units or unitary weightings [49, 25]. In case of dynamic loss weights that are not constant (e.g., UW), the weights do not cancel out fully due to the accumulation of gradient histories within Adam. Nonetheless, this has profound implications for loss weighting methods that are used in conjunction with Adam. For instance, when turning off the history within Adam (by setting $\beta_{1,2}=0$) and having a fixed backbone, all loss weighting methods, such as UW, random loss weighting and others, become equivalent to equal weighting.
Additional ablations to our previous experiments confirm the relevance of invariance in MTL (cf. LABEL:tab:res_nyu_segnet, LABEL:tab:res_nyu_deeplab, LABEL:tab:res_cs_segnet and LABEL:tab:res_cs_deeplab). First, we compare to signSGD+mom [3] which only updates on the sign of gradients and is therefore trivially scale-invariant in the heads. In a direct comparison with SGD+mom, we observe a superiority of signSGD for a majority of tested setups. Next, we applied task-specific Adam optimizers as in AdaTask [51] for a full loss-scale invariance and to allow an estimate of task-specific momentum and squared gradient accumulation. This is Pareto dominant over plain Adam+EW in almost all cases and significantly improves the $\Delta_{m}$ metric.
Conclusion: During MTL network training, we derive and measure a full loss-scale invariance for an optimal UW and a partial invariance for Adam. This partial invariance does not hold for SGD+mom and could explain, among other properties, the effectiveness of Adam in MTL as a baseline and its symbiotic behavior combination with other MTO methods. Thus, when comparing different loss weighting methods, it is crucial to be aware of the influence of the optimizer.

### v5

@@ Abstract
While multi-task learning (MTL) has gained significant attention in recent years, its underlying mechanisms remain poorly understood. Recent methods did not yield consistent performance improvements over single task learning (STL) baselines, underscoring the importance of gaining more profound insights about challenges specific to MTL. In our study, we investigate paradigms in MTL in the context of STL: First, the impact of the choice of optimizer has only been mildly investigated in MTL. We show the pivotal role of common STL tools such as the Adam optimizer in MTL empirically in various experiments. To further investigate Adam’s effectiveness, we theoretical derive a partial loss-scale invariance under mild assumptions. Second, the notion of gradient conflicts has often been phrased as a specific problem in MTL. We explore the role of gradient conflicts in MTL and compare it to STL. For angular gradient alignment we find no evidence that this is a unique problem in MTL. We emphasize differences in gradient magnitude as the main distinguishing factor. Overall, we find surprising similarities between STL and MTL suggesting to consider methods from both fields in a broader context.
@@ 1 Introduction
In a seminal study, Caruana [4] highlights both advantages and challenges in MTL. On the one hand, certain tasks can exhibit a symbiotic relationship, resulting in a mutual performance enhancement when trained together. On the other hand, conflicts between tasks can arise and decrease the performance when trained jointly, also known as negative transfer.
@@
The current understanding of MTL still lacks a deeper comprehension of its underlying mechanisms. To address this gap, our study aims to examine commonly held paradigms, such as the choice of optimizer, as well as the notion of gradient alignment and gradient magnitudes. Our contributions are:
@@
  • We examine the presumption of increased robustness on corrupted data as a result of MTL [24, 37]. We find light evidence that a higher number of tasks can result in improved transferability. Due to page limitations, we moved these results to App. A6, focusing the more compelling findings in the main text.
Overall, we provide a vast set of experiments and theoretical insights which contribute to a more comprehensive understanding of MTL in computer vision.
@@ 2 Related Work
Network architectures focus on the question of how features should be shared across tasks, e.g. [38, 34, 51, 36]. Multi-task optimization (MTO) aims to resolve imbalances and conflicts of tasks during MTL. Task affinities examine a grouping of tasks that should be learned together to benefit from the joint training [47, 12]. A general overview of recent works in MTL can be found in [48, 43]. Our work focuses on MTO, which we review more thoroughly in the following.
Gradient magnitude methods prevent the dominance of individual tasks by balancing them with task-specific weights. One line of works are loss-weighting methods. Here, weights are determined before any (task-wise) gradient computation and are used for a weighted aggregation of the tasks’ losses. These methods consider either the task uncertainty (UW) [21], rate of change of the losses (DWA, FAMO) [34, 30], the tasks’ difficulty (DTP) [14], validation performance by applying meta-learning (MOML, Auto-$\lambda$) [53, 33], or randomly chosen task weights (RLW) [28]. In line with these, the geometric mean of task losses has been used to handle the different convergence rates of the tasks [8]. An advantage of theses methods is their computational efficiency as the gradient needs to be computed only once for the aggregated loss. Alternatively, other methods consider the task-specific gradients directly, e.g., by normalizing them (GradNorm) [6] or propose a hybrid balancing between task-wise loss and gradient scaling (IMTL, DB-MTL) [32, 27]. Furthermore, there are several adaptions for the multiple-gradient descent algorithm (MGDA) [10], e.g. for applying it efficiently in deep learning setups [44] or by introducing a stochastic gradient correction [11]. Recently, task-wise gradient weights have been estimated by treating MTL as a bargaining problem (Nash-MTL) [40], or considering a stability criterion (Aligned-MTL) [45]. Crucially, all gradient magnitude methods consider scalar weightings of task-wise gradients within the backbone and/or heads. They do not modify the alignment of task-specific gradient vectors.
@@ 3 Problem Statement
Multi-task learning addresses the problem of learning a set of $T$ tasks simultaneously (see e.g. [48, 4]). It is noteworthy that this setup is occasionally also referred to as multi-label or multi-target learning [57] Importantly, this study does not incorporate multi-input data. We consider a supervised learning setup, use a shared backbone architecture, and learn all tasks together. Formally, given input data $\mathcal{X}$, the goal is to learn a function $f_{\bm{\theta}}(\mathbf{x})$ which maps a point $\mathbf{x}\in\mathcal{X}$ to each task label $y_{t}$ with $t=1,..,T$. The trainable parameters $\bm{\theta}=\{\phi,\psi_{1:T}\}$ consist of shared parameters $\phi$ and task-specific parameters $\psi_{t}$. Training a task $t$ is associated with the loss $\mathcal{L}_{t}(f_{\bm{\theta}}(x);\bm{\theta})$, e.g., a regression or classification loss. We denote respective gradients on the shared and task-specific parameters with $\mathbf{g}^{\phi}_{t}=\nabla_{\phi}\mathcal{L}_{t}$, and $\mathbf{g}^{\psi}_{t}=\nabla_{\psi}\mathcal{L}_{t}$. When training on multiple tasks, the shared parameters $\phi$ needs to be updated w.r.t. all task-wise gradients $\mathbf{g}^{\phi}_{t}$ which requires an appropriate aggregation. A simple solution is to uniformly sum up the task losses $\mathcal{L}=\sum_{t}\mathcal{L}_{t}$ which is referred to as Equal Weighting (EW). However, as tasks might be competing against each other, this can result in negative transfer and thus sub-optimal solutions. One way to deal with this difficulty is to adapt the magnitude of task-specific gradients. This can be achieved by weighting tasks during training, e.g., by scaling different losses $\mathcal{L}=\sum_{t}\alpha_{t}\mathcal{L}_{t}$, where $\alpha_{t}\geq 0$. Note that the $\alpha_{t}$ can change during training. Furthermore, the weighing can also be performed on gradient level to distinguish between shared and task-specific gradients. We refer to those approaches as gradient magnitude methods. Interestingly, the relationship between loss weights, network updates and learning rate also depends on the optimizer. We show a derivation for SGD and Adam in Section A1.2. Additionally to adapting the gradient magnitude, one can directly adapt the alignment of task-wise gradient vectors within the shared backbone $\tilde{\mathbf{g}}^{\phi}=\mathbf{h}(\mathbf{g_{1}}^{\phi},...,\mathbf{g_{T}}^{\phi})$.
@@ 4 Experiments and results
In this section we perform several experiments to gain a more profound understanding of multi-task learning (MTL) in computer vision by questioning common paradigms. We compare the impact of Adam and SGD in MTL in Sec. 4.1 and examine the process of gradient similarity in different settings in Sec. 4.2. Throughout this evaluation, we repeatedly make use of common setups, which we will specify as follows and in more detail in App. A3.
@@ 4.1 Effectiveness of Adam in multi-task learning
Examined paradigm: The impact of the choice of standard optimizer is often disregarded and varies across studies (overview in Tab. A1) when comparing MTL methods. For instance, Adam [22] was successfully used to show that random/constant weighting of tasks’ losses performs competitive compared to MTO methods [28, 25, 50]. In contrast, many methods proposing adaptive, task-specific weighting methods [32, 21] use stochastic gradient descent with momentum (SGD+mom). In recent works, the optimizer choice converged to Adam and a fixed learning rate schedule [34, 56, 31, 45] without a comparison to SGD+mom.
In this part of our study, we investigate the impact of Adam and SGD+mom in conjunction with common MTO methods. We identify the choice of optimizer as a crucial confounder in the experimental setup. Compared to SGD+mom, we find that the Adam optimizer itself is a quite effective baseline in MTL and can be regarded as a loss weighting method from a theoretical viewpoint.
@@ 4.1.1 Toy Task Experiment
Approach: We repeat the experiment of Liu et al. [31] using their original implementation but further test different learning rates and optimizers. They motivate their gradient alignment method CAGrad with a simple toy optimization problem in which their method reliably converges to the minimum of the average loss, while other MTO approaches would either get stuck (e.g., EW) or only converge to any point on the Pareto front (e.g., PCGrad [56], MGDA [44]).
Table 1: Maximum number of iterations for all seeds in the toy task experiment from [31] to reach the global minimum for varying MTO method, learning rate, and optimizer combination. EW+Adam often shows the fastest convergence to the global minimum. ’-’ denotes that not all seeds converged within 100k iterations. As reported in [31], PCGrad often only converges to a point on the Pareto Front. We highlight the best run for each learning rate over all MTO methods.
@@
Result: For higher learning rates with Adam optimizer, even equal weighting (EW) reaches the global optimum (cf. Fig. 1, e.g., EW+Adam, lr=0.05) and often converges even faster than dedicated MTO methods (Tab. 1). Note, original results were shown for learning rate 0.001 using Adam and were, thus, in favor of CAGrad. Results for additional learning rates are reported in Tab. A3.
Conclusion: The choice of optimizer appears to be more important on the success of the outcome of this experiment than the choice of MTO method, as Adam converges considerably faster and more reliably than GD. Also, tuning the learning rate is a relevant factor, however, especially in MTL with differently scaled losses, a single suitable learning rate for all tasks does often not exist.
@@ 4.1.2 Experiments on Cityscapes and NYUv2
We test the effectiveness of Adam and its role as a confounder in common MTL datasets for various MTO methods.
Approach: We compare Adam and SGD+mom in combination with any MTO method from equal weighting (EW), uncertainty weighting (UW) [21], random loss weighting (RLW) [28], PCGrad [56], CAGrad [31], IMTL [32] and Aligned-MTL [45], for which we used the implementation from  [29], as well as MTL-IO [41] and AdaTask [52]. We distinguish between any combination of dataset {Cityscapes [9], NYUv2 [39]} and network architecture {SegNet [1], DeepLabV3 [5]}. We run experiments for ten different initial learning rates from $[0.5,0.1,0.05,...,0.00001]$ and select the best one w.r.t. to the validation performance. More details are described in Sec. A3.2. As different models and parameter setups can show preference towards different tasks and metrics, we are interested in those models which are Pareto optimal (PO).
Table 2: Number of Pareto optimal (PO) experiments using either Adam or SGD+mom. as optimizer. Models trained with Adam are consistently more often on the Pareto front compared to those trained with SGD+mom. The number of Adam-based runs that are not dominated by any SGD-based run (PO w.r.t. SGD) is even higher, while the reverse does not apply.
@@
Results: We observe over all experimental setups that Adam performs favorably over SGD+mom (Tab. 2). This especially holds true for experiments on Cityscapes where the Pareto front for both network architectures only consists of Adam-based models. Moreover, an even larger number of Adam-based models is not dominated by any model trained with SGD+mom (PO w.r.t. SGD). For NYUv2, Adam still performs stronger but SGD+mom. also occasionally delivers a PO result. For the individual metrics, the predominance of Adam is further visualized in a parallel coordinate plot in Figs. 2 and A4. Bold lines indicate the overall Pareto optimal experiments (PO full).
In App. A4, we further report best $\Delta_{m}$ results for common MTO methods in combination with Adam or SGD+mom (App. A4, A4, A6 and A7). Again, Adam boosts the overall performance across methods. Furthermore, when comparing the ranking of MTO methods w.r.t. the $\Delta_{m}$ metric, we see that the order can change based on the choice of optimizer, e.g., for Cityscapes with SegNet the best method with Adam is UW but with SGD+mom it is CAGrad. This underlines the importance of the choice of optimizer as a confounder in the experimental setup. Noteworthy, EW with Adam yields Pareto optimal results in three of the four setups (cf. Tab. A2) and is not dominated by any specialized MTO method trained in combination with SGD+mom for all dataset and network combinations. This supports claims questioning the effectiveness of specific MTO methods [25, 50]. Nonetheless, looking at the $\Delta_{m}$ metric and individual metrics, we see that sometimes with a small relative performance drop on one metric, significant gains on another metric can be achieved (e.g., Cityscapes+sem.seg. and depth for UW vs EW).
@@ 4.1.3 The reasonable effectiveness of Adam in the context of uncertainty weighting
We show that Adam’s mechanism to estimate a parameter-specific learning rate is partially loss-scale invariant and hypothesize that this could contribute to Adams effectiveness in MTL. We demonstrate this partial invariance theoretically and empirically. Furthermore, a full loss-scale invariance can also be shown under mild assumptions for UW [21], which is among the most prevalent loss weighting method in the literature, and related similar variant [27].
The loss-scale invariance of UW can be shown by assuming an optimal solution for the $\sigma$ values similar to [23]. This assumption is mild as this is a 1-dimensional convex optimization problem for each $\sigma$. The invariance can be demonstrated by inserting the analytical solution starting from UW. For example, assuming a Laplacian distribution (this can be shown for other distributions as well), we have
@@
where $sg$ is the stop-gradient operator and $c$ is a constant that can be omitted during optimization. Given this, we directly see the invariance w.r.t. loss-scalings. For instance, with $\mathcal{L}_{1}\rightarrow\alpha_{1}\mathcal{L}_{1}$ and $\mathcal{L}_{2}\rightarrow\alpha_{2}\mathcal{L}_{2}$, the derivative of the total loss $\mathcal{L}$ remains unchanged. As this invariance is shown on the loss-level, it holds for all gradient updates w.r.t. the head and backbone. Intuitively, this could explain why UW performs strongly in the context of various loss scalings such as measuring depth in centimeters or meters. Further details, are in App. A1.
@@
Thus, for the network heads, we see a similar effect as for optimal UW that different scalings do not impact the network update. However, this does not hold for the backbone. The full derivation is shown in App. A1. We confirm empirically in a handcrafted loss-scaling experiment in App. A2 and Figs. A1 and A2 that SGD does not offer any scaling invariance, whereas Adam involves the invariance property for the heads. The optimal UW demonstrates a scaling invariance for the heads and the backbone.
We would like to note that our derivation for Adam is only valid for constant $\alpha_{t}$, e.g., measuring depth in different units or unitary weightings [50, 25]. In case of dynamic loss weights that are not constant (e.g., UW), the weights do not cancel out fully due to the accumulation of gradient histories within Adam. Nonetheless, this has profound implications for loss weighting methods that are used in conjunction with Adam. For instance, when turning off the history within Adam (by setting $\beta_{1,2}=0$) and having a fixed backbone, all loss weighting methods, such as UW, RLW, and others, become equivalent to equal weighting.
Additional ablations to our previous experiments suppport the relevance of invariance in MTL (cf. Tabs. A6, A7, A4 and A4). First, we compare to signSGD+mom [3] which only updates on the sign of gradients and is therefore trivially scale-invariant in the heads. In a direct comparison with SGD+mom, we observe a superiority of signSGD for a majority of tested setups. Next, we applied task-specific Adam optimizers as in AdaTask [52] for a full loss-scale invariance and to allow an estimate of task-specific momentum and squared gradient accumulation. This is Pareto dominant over plain Adam+EW in almost all cases and significantly improves the $\Delta_{m}$ metric.
Conclusion: In the context of MTL, we derive and measure a full loss-scale invariance for an optimal UW and a partial invariance for Adam. This partial invariance does not hold for SGD+mom and could explain, among other properties, the effectiveness of Adam in MTL. Furthermore, when comparing different loss weighting methods, it is crucial to be aware of the influence of the optimizer.